# “I think this is fair”: Uncovering the Complexities of Stakeholder Decision-Making in AI Fairness Assessment

### v1

@@ 5.2. Patterns in AI Fairness Decision-making
Tailoring metrics to features: Unlike the “one-size-fits-all” approach commonly adopted by AI experts, our findings uniquely observed that stakeholders selected different fairness metrics depending on the specific feature under consideration. Notably, this pattern has been largely absent in prior research (Cheng et al., 2021), which often offers limited metric choices or discusses metric preferences in isolation from specific features. In child maltreatment risk prediction, where, within a constrained set of protected features and fairness metrics, participants tended to maintain a consistent metric preference across features (Cheng et al., 2021). However, this pattern may generalize to domains where features carry distinct fairness risks and stakeholders engage with a wide range of fairness definitions.
Favoring custom metrics: Our participants showed strong interest in developing custom metrics tailored to their needs, an aspect that has not been examined in prior research (Nakao et al., 2023). This divergence may stem from stakeholders’ reliance on lived experiences and contextual knowledge, which resist abstraction into standardized formulas. It also surfaces a tension between the expert-driven pursuit of comparability (favoring fixed metrics) and the stakeholder-driven pursuit of legitimacy (favoring flexible, situated definitions). This tension is unlikely to be confined to the loan scenario and may recur across other high-stakes domains where fairness cannot be meaningfully captured by one-size-fits-all solutions.
@@
Setting stricter thresholds: We found that some stakeholders initially aspired to perfect fairness outcomes based on their chosen feature–metric pairs but introduced tolerance when considering practical feasibility. Importantly, tolerance for unfairness decreased as their perceived harm increased. This pattern echoes prior research (Luo et al., 2025) and suggests a generalizable logic in how stakeholders set fairness thresholds.
Shifts in fairness priority: Shifts in feature rankings suggest that stakeholder fairness judgments evolve with exposure to information and trade-offs, rather than reflecting fixed preferences. This dynamism also challenges one-shot, expert-centric evaluations.
5.3. Implications for Fairness Assessment Tool Design
Building on the observed assessment patterns and user feedback (Section 4.3) from our current prototype, we outline design implications in Table 2 for future fairness assessment tools to better support stakeholder involvement, extending the design requirements summarized in Table 1 (Nakao et al., 2023).
@@ 5.3. Implications for Fairness Assessment Tool Design
Table 2. Design Implications for Tools Supporting Stakeholder Fairness Assessments
Category Design Implications Description
Feature Selection
(Whom to Protect)
Explain features Provide plain-language explanations for features used in predictive AI models.
@@
Support diverse feature selection Allow users to add or remove both protected and non-protected features, rather than defaulting to protected features only.
@@
Enable feature ranking Allow users to rank features by perceived priority.
Metric Selection
(How to Protect)
@@
Offer diverse metrics Support both group- and individual-level fairness metrics, including both widely used and less common ones, to avoid defaulting only to expert preferences.
@@
Support custom metrics Allow users to combine multiple metrics, specify their logical relationships (e.g., and/or), assign weights, or provide free-text input to flexibly express desired fairness metric.
@@
Dynamic assessment Allow continuous modification of features, metrics, and rankings during the fairness assessment process.
Threshold Setting
(To What Extent)
Thresholds per feature-metric pair Allow threshold input as either a single value or a range.
5.4. Scaling Up Stakeholder Involvement
In real-world practice, stakeholder involvement often requires engaging a significantly larger and more diverse population than was feasible in our study. Our current approach, where researchers interact with participants one-on-one in person, is neither scalable nor efficient due to substantial time and resource demands. Consequently, real-world deployment calls for systems that are accessible online and support simultaneous, independent use by multiple stakeholders. This necessitates thoughtful system design that communicates tasks and fairness metrics clearly, while minimizing cognitive load and improving user experience (Luo et al., 2025). However, the design and real-world deployment of such tools remain challenging in supporting users’ understanding of AI fairness (Kallina and Singh, 2024), and this area remains underexplored in existing research (Nakao et al., 2023).
@@ 6. Conclusion
We investigated the complexity of stakeholders’ fairness assessments by examining not only who they want to protect (feature selection and prioritization), but also how they want to protect them (feature–metric pairing), and what level of unfairness they consider acceptable (threshold setting). To support this process, we developed a prototype system and conducted a qualitative study in a credit rating scenario with 30 participants. Our findings show that non-experts can engage in these expert-like tasks through the UI, yet do so in ways that are more context-sensitive, diverse, and stringent than expert defaults:
@@
  • Participants set stricter fairness thresholds than conventional legal or technical practice, reflecting higher demands for fairness but this is also influenced by personal preferences.
@@ Acknowledgements.
We gratefully acknowledge the funding provided by the University of Glasgow and Fujitsu Limited. We thank all participants for their contributions.
@@ Appendix A Appendix
A.1. Participant Information
This is the demographic information of all 30 participants, collected from the background questionnaire. Participants ranged in age from 21 to 76, with 19 aged 20–30 years, 8 aged 30–40 years, 2 aged 40–50 years, and 1 over 50. Educational backgrounds ranged from secondary school (n=1) to bachelor’s (n=20), master’s (n=6), and doctoral degrees (n=3). The sample included 13 male and 17 female participants, with 14 reporting prior loan application experience and 16 without such experience.
Table 3. Participant Information
@@ 4 Yes 24 Female Bachelor’s Degree Computer and IT Indian
5 No 43 Male Bachelor’s Degree Political and Social Studies Pakistani
@@ 7 Yes 24 Female Bachelor’s Degree Healthcare and Medical Iranian
8 Yes 33 Female Doctoral Degree Engineering Chinese
9 No 34 Female Doctoral Degree Education Indian
@@ 12 No 24 Male Bachelor’s Degree Computer and IT Chinese
13 Yes 29 Female Bachelor’s Degree Healthcare and Medical Chinese
14 Yes 32 Male Bachelor’s Degree Business and Finance Indian
@@ 24 No 23 Male Bachelor’s Degree Business and Finance Chinese
25 Yes 37 Female Bachelor’s Degree Chemical Biology Iranian
@@ 30 No 33 Male Bachelor’s Degree Education Bangladeshi
A.2. Post Questionnaires
@@ 1 Strongly disagree    2 Disagree    3 Neutral    4 Agree   5 Strongly Agree
A.3. Prototype System - Detailed Description
A.3.1. Supplementary Material for Task 1: Feature Selection & Ranking
Participants were able to click on “Causal Graph: Training Data” to check the training data’s casual graph. Each node’s (feature) influence on the other nodes was shown by arrows, with the thickness of these arrows indicating the strength of the relationship. Orange nodes indicated protected features, whereas the dark blue node represented the bank customer’s Rated Credit (Ground truth label). Participants were able to click on each node to highlight the influence other nodes had on the selected node and the influence the selected node had on the other nodes via the direction of the arrows. Similarly, Participants could click on “Causal Graph: Test Data” and interacted with the test data causal graph with the dark blue node indicating AI’s Predicted Credit. In addition to exploring data bias, participants could click on “Feature Importance” to examine potential model biases. Our prototype displayed a bar chart with feature importance weights on the X-axis and features on the Y-axis, sorted by importance. Green bars (weights ¿ 0) indicated positive impacts, red bars (weights ¡ 0) negative impacts. Participants could also click on “Data Distribution” to view multiple pie charts, each representing the distribution of a feature in the training data.
@@ A.3.1. Supplementary Material for Task 1: Feature Selection & Ranking
A.3.2. Supplementary Material for Task 2: Mapping Features to Metrics with Thresholds
Within the “Metric Exploration” component, participants were able to start by clicking “GROUP FAIRNESS” or “INDIVIDUAL FAIRNESS” to choose the fairness category to explore. When exploring group fairness, participants could select a feature from the dropdown menu, arranged by their own feature ranking. Once selected, participants could view the performance of the chosen feature in terms of group fairness. Participants could click on a specific metric, the explanation content would be updated correspondingly, displaying a plain-language description. After clicking “Conditional Statistical Parity”, participants could customize the condition feature by clicking the dropdown menu, then view results and instance-level visualizations based on the selected condition and feature (Figure  5).
When exploring individual fairness, participants were able to click on “COUNTERFACTUAL FAIRNESS” and chose protected and sensitive features via radio buttons (Figure  6). This displayed visualizations and explanations of counterfactual fairness, highlighting examples of individuals who violate counterfactual fairness. Participants could also select any non-protected feature using the dropdown in our prototype system, indicating their interest in the corresponding counterfactual fairness for that feature. For consistency, following FEE’s design, participants could click “CONSISTENCY” to see metric results and the predictions of each instance’s neighbors, with red dots indicating good predictions and green for the opposite.
@@ A.3.2. Supplementary Material for Task 2: Mapping Features to Metrics with Thresholds
A.3.3. Supplementary Material for Task 3: Fairness Re-Ranking
@@
A.4. Fairness Metrics and Explanations
As referenced in (Luo et al., 2025), the table below provides definitions and calculations for all eight metrics used in our study.

### v2

@@ 5.2.4. Tailoring metrics to features
Unlike the “one-size-fits-all” approach commonly adopted by AI experts (de Castro Vieira et al., 2025; Hort et al., 2024), we observed that a relatively small number of stakeholders used the same metric across all features, while most applied different fairness metrics depending on the specific feature under consideration. Notably, this pattern has not clearly emerged in prior research (Cheng et al., 2021), which often offers limited metric choices or discusses metric preferences in isolation from specific features. In child maltreatment risk prediction, where, within a relatively smaller set of protected features and fairness metrics, participants tended to maintain a consistent metric preference across features (Cheng et al., 2021). This difference further reinforces the need for fairness-assessment tools that not only support context-specific reasoning but also surface feature-specific fairness needs.
5.2.5. Threshold setting as a reflection of high fairness expectations
Considering that fairness metrics differ not only across categories (e.g., group vs. individual fairness) but also within them (e.g., Counterfactual Fairness vs. Consistency) (Mehrabi et al., 2021), we carefully defined thresholds as the maximum allowable deviation from the optimal fairness outcome of the current metric, for participants, they were maximum unfairness participants were willing to tolerate (0–100%) to reduce cognitive load while keeping the concept clear and consistent. Although no unanimous agreement on specific fairness threshold values emerged due to personal preferences, participants consistently showed strict settings, revealing high fairness expectations and low tolerance for unfairness.
@@ 5.2.5. Threshold setting as a reflection of high fairness expectations
We also found that some participants initially aspired to ‘perfect’ fairness outcomes for their chosen feature–metric pairs but introduced tolerance when considering practical feasibility. Importantly, tolerance for unfairness decreased as their perceived harm increased. This pattern echoes prior research (Luo et al., 2025) and therefore suggests that stakeholders may set thresholds by balancing high fairness expectations with practical feasibility. Despite this, AI expert practice typically allows for larger deviations (10% or 20%), often due to legal, operational, or organizational constraints (Feldman et al., 2015; Gupta et al., 2021), than the thresholds preferred by participants (mostly below 6%). This gap suggests that AI experts may face greater practical demands during bias mitigation when attempting to align their models with stakeholder expectations.
However, current threshold definitions (i.e., the maximum allowable deviation from a metric’s ideal fairness outcome, expressed as 0–100%) might lack comparability across heterogeneous fairness metrics. Normalization schemes could help when a metric’s value range is clearly defined. For instance, normalizing all fairness metrics to a common scale (e.g., 0-1) would make their threshold values directly comparable and help AI experts to better analyze which metrics, or which properties of those metrics, influence stakeholders to set more stringent or more permissive fairness thresholds. Even after normalization, fairness metrics may differ in whether larger (i.e., close to 1) or smaller values (i.e., close to 0) indicate greater fairness. To make thresholds comparable across all metrics, we could adopt a single convention in which lower values always represent better fairness and values below a chosen threshold are considered acceptable. Metrics that originally follow the opposite direction are transformed by reversing their scale: a normalized value $x$ used for thresholding is replaced with 1x1-x so that, for example, a threshold value originally considered fair at $\geq 0.9$ becomes an equivalent fair value at $\leq 0.1$. However, a metric’s value range is not always straightforward to determine, especially when it is unbounded and lacks a known maximum, for example, in the case of stakeholder-defined custom fairness metrics. This can limit the applicability of many normalization schemes and introduce additional effort for AI experts.
5.2.6. Fairness priorities are not always static.
We observed shifts in participants’ fairness priorities as they engaged in assessing fairness. Such dynamics may be driven by considerations of contextual relevance and trade-offs as participants are exposed to more fairness- and context-related information. This dynamism also resonates with recent calls from AI experts, realizing that fairness assessment is not a static setting (Tang et al., 2023), though it remains underexplored in stakeholder fairness assessments. This warrants further investigation and should be carefully considered in the design of long-term fairness assessment tools.
@@
5.3. Design Implications for Tools Supporting and Scaling Stakeholder Fairness Assessments
In real-world practice, stakeholder involvement often requires engaging a significantly larger and more diverse population than was feasible in our study. Our current approach, where researchers interact with participants one-on-one in person, is neither scalable nor efficient due to substantial time and resource demands. Consequently, real-world deployment calls for tools that are accessible online and support independent use by multiple stakeholders. However, the latter remains challenging and underexplored (Kallina and Singh, 2024; Luo et al., 2025; Nakao et al., 2023; Tahaei et al., 2023).
Recent work highlights the need for fairness assessment tools that help stakeholders move from “aspirational fairness” (“what (fairness) should be”) to a more technical “measurement view” (“how is it working or performing”) (Müller et al., 2025), and that present fairness concepts to the general public with minimal prior knowledge (Tahaei et al., 2023). Building on these insights, we suggest that on one hand, tools should support stakeholders in assessing AI fairness in ways that accurately reflect their needs (functional); on the other hand, tools should also communicate fairness concepts clearly and support independent use (informational), thereby enabling scaling.
We outline the first iteration of design implications in Table 2 based on our prototype that speaks to both functional and informational needs, supporting and scaling lay stakeholders in assessing AI outcome fairness. These implications are grounded in the stakeholder assessment patterns we observed (Section 5.2) and the challenges participants reported during the study (Section 4.4).
Table 2. Design Implications for Tools Supporting and Scaling Lay Stakeholder Fairness Assessments
Category Use Design Implications Description
@@ 5.3. Design Implications for Tools Supporting and Scaling Stakeholder Fairness Assessments
Feature Selection (Whom to Protect) Functional Support diverse feature selection Allow stakeholders to add or remove both protected and non-protected features, rather than defaulting to protected features only.
@@
Enable feature ranking Allow stakeholders to rank features by perceived priority.
Informational Provide clear and accessible feature descriptions Provide descriptions for all features to help stakeholders, including those unfamiliar with the application domain, clearly understand each feature.
Support navigation within large feature sets For high-dimensional datasets, avoid unstructured feature lists, for example by presenting related features in coherent orders, and surface manageable subsets when needed while maintaining scenario completeness and avoiding expert-driven selection bias (Cheng et al., 2021; Grgić-Hlača et al., 2018; Zhang et al., 2023).
Scaffold user reflection on fairness Help elicit fairness concerns for those unsure where to begin reasoning about fairness without leading them, for example through reflection on experiences of unfairness (Smith et al., 2024), or historical bias.
@@
Metric Selection (How to Protect) Functional Offer diverse metrics Support both group- and individual-level fairness metrics, including both widely used and less common ones, to avoid defaulting only to expert preferences.
@@
Support custom metrics Allow stakeholders to combine multiple metrics, specify their logical relationships (e.g., and/or), assign weights, indicate metric priorities, or provide free-text input to flexibly express desired fairness metric.
@@
Allow dynamic assessment Allow continuous modification of features, metrics, and rankings during the fairness assessment process.
Informational Offer multi-modal explanations while minimizing information load. Offer multi-modal explanations for each fairness metric, such as lay-language descriptions, visual examples of when a metric is satisfied or violated, result visualizations, etc., and surface them progressively to reduce information load and address the challenge of knowing where to focus.
Offer a comparative view of multiple fairness metrics Highlight key characteristics of multiple metrics within the metric comparison view, for example their strengths, limitations, and distinctions, allowing stakeholders to compare them easily without relying on memory or switching between metrics.
Threshold Setting (To What Extent) Functional Thresholds per feature-metric pair Allow threshold input as either a single value or a range.
Informational Explain fairness thresholds Describe what the fairness threshold represents and show its valid range for the selected feature–metric pair to support stakeholders’ understanding.
Provide real-time fairness assessment feedback Show instantly whether the model is fair under the current threshold to validate stakeholders’ understanding and ensure the threshold aligns with their intent (Luo et al., 2025).
@@ 6. Conclusion
We investigated the complexity of stakeholders’ fairness assessments for outcome fairness by examining not only who they want to protect (feature selection and prioritization), but also how they want to protect them (feature–metric pairing), and what level of unfairness they consider acceptable (threshold setting). To support this process, we developed a prototype system and conducted a qualitative study in a credit rating scenario with 26 participants. Our findings show that lay stakeholders can express fairness assessments, and do so in complex ways with fairness considerations that are contextually grounded, broad in scope, dynamic, customized, and held to high expectations:
@@
  • Participants tended to set fairness thresholds that were strict yet varied, reflecting high expectations for AI fairness while also being influenced by personal preferences.
@@ Acknowledgements.
We gratefully acknowledge the funding provided by the University of Glasgow and Fujitsu Limited. This work was also partially supported by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from Responsible Ai UK (KP0011). We thank all participants for their contributions.
@@ Appendix A Appendix
A.1. German Credit Dataset: Features and Feature Descriptions
Table 3 shows 20 features with their descriptions from the German Credit Dataset that participants could refer to.
Table 3. German Credit Dataset: Features with their Descriptions
Feature Name Feature Description Type Example Value
Duration Duration of the credit in months Numerical 30 (Months)
Credit Amount Credit amount asked by the applicant Numerical 5234 (DM)
Installment Rate Installment rate in percentage of disposable income Numerical 4
Residence Length Length of time (in years) the applicant has lived in the present residence Numerical 1
Existing Credits Number of existing credits at this bank Numerical 1
Dependents Number of people being liable to provide maintenance for Numerical 1
Age The age of the applicant in years Numerical 28
Gender The gender of the applicant Categorical Male
Checking Account Status/balance of checking account at this bank Categorical 0-200 DM
Credit History Past credit history of applicant at this bank Categorical Other Existing Credits
Purpose The applicant’s purpose for the loan (e.g., car, education) Categorical Car (New)
Savings Savings accounts/bonds at this bank Categorical ¡100 DM
Employment Present employment since Categorical Unemployed
Debtors Other debtors/guarantors present Categorical None
Property Properties that applicant has Categorical Car/Other
Installment Plans Other installment plans the applicant is paying Categorical None
Housing Housing (rent, own, for free) Categorical Own
Job Current job information (unemployed, (un)skilled, management) Categorical Management/Officer
Telephone Is there any telephone registered for this applicant? Binary Yes
Foreign Worker Is applicant foreign worker? Binary Yes
A.2. Participant Information
Table 4. Participant Information: Participants are grouped by fairness perspective (outcome fairness vs. other perspectives) rather than listed by PID to support easier cross-referencing with results; PID 9, 22, and 25 were also placed in a separate set, since part of their data did not reflect the outcome fairness perspective (Section 4.1).
Participants with Outcome Fairness Perspective
@@ 30 No 33 Male Bachelor’s Degree Education Bangladeshi
9 No 34 Female Doctoral Degree Education Indian
22 Yes 35 Male Master’s Degree Education British
25 Yes 37 Female Bachelor’s Degree Chemical Biology Iranian
Participants with Other Fairness Perspectives
5 No 43 Male Bachelor’s Degree Political and Social Studies Pakistani
8 Yes 33 Female Doctoral Degree Engineering Chinese
13 Yes 29 Female Bachelor’s Degree Healthcare and Medical Chinese
14 Yes 32 Male Bachelor’s Degree Business and Finance Indian
Table 5. Participant Information Summary: Participants with outcome fairness perspective were included in the analysis, whereas others were excluded.
Summary of Participants with Outcome Fairness Perspective (N=26)
Loan Application Experience Age Gender Education Profession Nationality
Yes (11) aged 20–30 years (18) Male (11) Secondary school diploma or equivalent (1) Biology (1) Saudi Arabian (1)
No (15) aged 30–40 years (6) Female (15) Bachelor’s Degree (17) Healthcare and Medical (6) Nigerian (1)
aged 40–50 years (1) Master’s Degree (6) Construction and Manufacturing (1) Chinese (7)
over 50 years (1) Doctoral Degree (2) Computer and IT (3) Indian (7)
Education (6) Iranian (3)
Engineering (2) Egyptian (1)
Business and Finance (4) Irish (1)
Chemical Biology (1) British (3)
Physics (1) Bangladeshi (2)
Summary of Participants with Other Fairness Perspective (N=4)
Loan Application Experience Age Gender Education Profession Nationality
Yes (3) aged 20–30 years (1) Male (2) Bachelor’s Degree (3) Political and Social Studies (1) Pakistani (1)
No (1) aged 30–40 years (2) Female (2) Doctoral Degree (1) Engineering (1) Chinese (2)
aged 40–50 years (1) Healthcare and Medical (1) Indian (1)
Business and Finance (1)
A.3. Post Questionnaires
@@ 1 Strongly disagree    2 Disagree    3 Neutral    4 Agree   5 Strongly Agree
A.4. Prototype System - Detailed Description
A.4.1. Supplementary Material for Task 1: Feature Selection & Ranking
Participants were able to click on “Causal Graph: Training Data” to check the training data’s casual graph. Each node’s (feature) influence on the other nodes was shown by arrows, with the thickness of these arrows indicating the strength of the relationship. Orange nodes indicated protected features, whereas the dark blue node represented the applicant’s Rated Credit (Ground truth label). Participants were able to click on each node to highlight the influence other nodes had on the selected node and the influence the selected node had on the other nodes via the direction of the arrows. Similarly, Participants could click on “Causal Graph: Test Data” and interacted with the test data causal graph with the dark blue node indicating AI’s Predicted Credit. In addition to exploring data bias, participants could click on “Feature Importance” to examine potential model biases. Our prototype displayed a bar chart with feature importance weights on the X-axis and features on the Y-axis, sorted by importance. Green bars (weights ¿ 0) indicated positive impacts, red bars (weights ¡ 0) negative impacts. Participants could also click on “Data Distribution” to view multiple pie charts, each representing the distribution of a feature in the training data.
@@ A.4.1. Supplementary Material for Task 1: Feature Selection & Ranking
Explanation of the Conditional Statistical Parity metric, combining textual description with an intuitive instance-level visualization.
A.4.2. Supplementary Material for Task 2: Mapping Features to Metrics with Thresholds
Within the “Metric Exploration” component, participants were able to start by clicking “GROUP FAIRNESS” or “INDIVIDUAL FAIRNESS” to choose the fairness category to explore. When exploring group fairness, participants could select a feature from the dropdown menu, arranged by their own feature ranking. Once selected, participants could view the performance of the chosen feature in terms of group fairness. Participants could click on a specific metric, the explanation content would be updated correspondingly, displaying a plain-language description. After clicking “Conditional Statistical Parity”, participants could customize the condition feature by clicking the dropdown menu, then view results and instance-level visualizations based on the selected condition and feature (Figure  4).
When exploring individual fairness, participants were able to click on “COUNTERFACTUAL FAIRNESS” and chose protected and sensitive features via radio buttons (Figure  5). This displayed visualizations and explanations of counterfactual fairness, highlighting examples of individuals who violate counterfactual fairness. Participants could also select any non-protected feature using the dropdown in our prototype system, indicating their interest in the corresponding counterfactual fairness for that feature. For consistency, following FEE’s design, participants could click “CONSISTENCY” to see metric results and the predictions of each instance’s neighbors, with red dots indicating good predictions and green for the opposite.
@@ A.4.2. Supplementary Material for Task 2: Mapping Features to Metrics with Thresholds
Explanation of the Counterfactual Fairness metric, combining textual description with a supporting visualization to illustrate how outcomes differ under counterfactual conditions.
A.4.3. Supplementary Material for Task 3: Fairness Re-Ranking
@@
Prototype system module for Task 3. The top section presents a textual task description, and the bottom section lists feature–metric pairs that participants can rank to explore fairness trade-offs.
A.5. Fairness Metrics and Explanations
As referenced in (Luo et al., 2025), Table 6 provides definitions and calculations for all eight metrics used in our study.